OpenStudy (anonymous):

Find an equation in point slope- form for the line having the specified slope and containing the point indicated m=4, (3,8)

OpenStudy (anonymous):

y=4x+c Sub in the point 8=4*3+c 8=12+c c=-4 therefore final equation: y=4x-4

OpenStudy (anonymous):

Thanks a lot I actually did not have much clue on this one